OpenStudy (anonymous):

show that the function f(x)=(x-3)^(1/2) (2x-5) has derivative f'(x)= (6x-17) / 2(X-3)^(1/2)

OpenStudy (anonymous):

how far did you get?

OpenStudy (anonymous):

remember to find a common denominator after you take the derivative...

OpenStudy (anonymous):

i got f'(x)= \[1/2(x-3)^{-1/2}(2x-5)+(x-3)^{1/2}(2)\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

now find a CD (2*sqrt(x-3))

OpenStudy (anonymous):

do you see where this is going?

OpenStudy (anonymous):

one is negative and one is positive though?

OpenStudy (anonymous):

\[\frac{ 2x-5 }{ 2\sqrt{x-3} } + \frac{ 2\sqrt{x-3} }{ 1 }*\frac{ 2\sqrt{x-3} }{ 2\sqrt{x-3} }\]

OpenStudy (anonymous):

algebra flashback occurring ...

OpenStudy (anonymous):

ohhhhh right!